OpenStudy (anonymous):

The composition functions f?g and g?f are equal to each other for any f and g. This isnt true, right? My only question is what is a "composite function" example: f=x-1 g=x^2 f(g(x))=x^2-1 g(f(x))=x^2 -2x +1

OpenStudy (imstuck):

A composite function is a function made up of 2 functions. You "do" g(x) to f(x) here to find the result.

OpenStudy (anonymous):

g(f(x)) should be (x-1)^2 instead

OpenStudy (anonymous):

(x-1)(x-1) x^2-x x^2-x-x x^2-x-x+1 x^2 -2x +1 i wrote it correctly i believe

OpenStudy (anonymous):

soooo this is false?

OpenStudy (anonymous):

the statement " f?g and g?f are equal to each other for any f and g." is false

OpenStudy (anonymous):

thanks:)

OpenStudy (perl):

f=x-1 g=x^2 f(g(x))=x^2-1 g(f(x))=(x-1)^2 = x^2 -2x +1

OpenStudy (anonymous):

yes your x^2 -2x +1 is correct too

OpenStudy (perl):

the reason why it is false it is because it says 'for any f,g f compose g = g compose f' . This is not true generally (it is true when f and g are inverses of each other) . So all you have to do is find a counterexample. and you found a counterexample

OpenStudy (perl):

To falsify a general statement such as above , it is sufficient to find one counterexample.

OpenStudy (anonymous):

@perl do you know which function is the reflection on y=x ?

OpenStudy (anonymous):

inverse function is reflection on y=x right ?

OpenStudy (perl):

the inverse function , i think

OpenStudy (anonymous):

ok thanks!

OpenStudy (perl):

it is not necessarily a function, to flip x and y . sometimes we just call it the 'inverse relation'

OpenStudy (perl):

for example if you reflect y = x^2 about the line y = x , you don't get a function

OpenStudy (anonymous):

@perl can you take a look at my product moment coefficient qn ?